OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2-12x+7=0

OpenStudy (solomonzelman):

add 29 to both sides.

OpenStudy (solomonzelman):

can you add 29 to both sides?

OpenStudy (anonymous):

which sides?

OpenStudy (solomonzelman):

Both. \[x^2-12+7+29=0+29\]

OpenStudy (anonymous):

and why 29?

OpenStudy (solomonzelman):

To make the left side into perfect square because you have to have (-6) (-6) which equals 36.

hero (hero):

x^2 - 12x + 7 = 0 x^2 - 12x = -7 x^2 - 12x + (-12/2)^2 = (-12/2)^2 - 7 x^2 - 12x + 36 = 36 - 7 (x - 6)^2 = 29 Take square root of both sides to finish solving for x

OpenStudy (anonymous):

Oh. Okay. I still don't quite get it, but lets move on. What shall I do next? :D

OpenStudy (anonymous):

Uhh... which one is right?

OpenStudy (gorv):

x^2-2*x*6+7=0 x^2-2*x*6+36-36+7=0 (x^2-2*x*6+36)-29=0 (x-6)^2-29=0 (x-6)^2=29 x-6=29^1/2 x=(29)^1/2+6

OpenStudy (anonymous):

Are all of these right, just in different ways?

hero (hero):

You add 36 to both sides, not 29

hero (hero):

But 36 - 7 = 29

OpenStudy (solomonzelman):

@Hero, no you add 29, because that makes the left side into perfect square. @Emily778, I know what you are confused about, lets learn how to complete the square.

OpenStudy (anonymous):

I'll just do this x^2 - 12x + 7 = 0 x^2 - 12x = -7 x^2 - 12x + (-12/2)^2 = (-12/2)^2 - 7 x^2 - 12x + 36 = 36 - 7 (x - 6)^2 = 29

hero (hero):

@SolomanZelman, if that's true, then explain how I got my result

OpenStudy (gorv):

actually in (a+b)^2=a^2+b^2+2ab here 2ab is 12x=2*x*6 so u need to make b^2 that is 36

hero (hero):

I know exactly how the process works and what to add to both sides.

OpenStudy (solomonzelman):

@Emily778, do you get how I came up with 29? lets do a different example. GUYS PLEASE, LETS TEACH HER, NOT DO HER WORK> She needs to understand.

OpenStudy (gorv):

@hero if u will add 29 both side than on x^2-12x+7+29=29 x^2-12x+36=29

hero (hero):

She'll only understand if those teaching know what they are talking about

OpenStudy (solomonzelman):

STOP ARGUING!

OpenStudy (solomonzelman):

YES, add 29

OpenStudy (gorv):

@hero plzz plzzzz plzzzz for god sake use little of ur mind

OpenStudy (solomonzelman):

@Emily778, do you get how I came up with 29? Do you want to learn how to complete the square?

hero (hero):

I think @Emily778 already has it figured out lol

OpenStudy (solomonzelman):

MODERATOR, please don't ruin the thread, and confuse teh asker, lets teach her!

OpenStudy (anonymous):

@Hero What do I do after this? Or is it already done? x^2 - 12x + 7 = 0 x^2 - 12x = -7 x^2 - 12x + (-12/2)^2 = (-12/2)^2 - 7 x^2 - 12x + 36 = 36 - 7 (x - 6)^2 = 29

OpenStudy (solomonzelman):

Sorry If I was offensive, @Hero .

hero (hero):

Finish solving for x @Emily778

hero (hero):

Take the square root of both sides

OpenStudy (anonymous):

Can you show me how? I am so confused:(

OpenStudy (solomonzelman):

@Emily778 PLEASE read this CAREFULLY. Do you know how to come up with the number that you have to add to both sides to ake the left side into a perfect square?

OpenStudy (solomonzelman):

to make the left side...typo

hero (hero):